Use the lever system equation. F1x=F2(d-x)
An adult and child are on a seesaw that is 15 ft long. The adult weighs 192 lb. And the child weighs 48 lb. How many feet from the child must the fulcrum be placed so that the seesaw is balances?

Answer by stanbon(75887)

Use the lever system equation. F1x=F2(d-x)
An adult and child are on a seesaw that is 15 ft long. The adult weighs 192 lb. And the child weighs 48 lb. How many feet from the child must the fulcrum be placed so that the seesaw is balances?
------
192x = 48(15-x)
-------
192x = 48*15 - 48x
-------
240x = 720
x = 3 ft (distance from the adult)
15-3 = 12 ft (distance from the child)
-------------
Cheers,
Stan H.
-------------